# 3M-Health: Multimodal Multi-Teacher Knowledge Distillation for Mental Health Detection

### v2

@@ Abstract.
The significance of mental health classification is paramount in contemporary society, where digital platforms serve as crucial sources for monitoring individuals’ well-being. However, existing social media mental health datasets primarily consist of text-only samples, potentially limiting the efficacy of models trained on such data. Recognising that humans utilise cross-modal information to comprehend complex situations or issues, we present a novel approach to address the limitations of current methodologies. In this work, we introduce a Multimodal and Multi-Teacher Knowledge Distillation model for Mental Health Classification, leveraging insights from cross-modal human understanding. Unlike conventional approaches that often rely on simple concatenation to integrate diverse features, our model addresses the challenge of appropriately representing inputs of varying natures (e.g., texts and sounds). To mitigate the computational complexity associated with integrating all features into a single model, we employ a multimodal and multi-teacher architecture. By distributing the learning process across multiple teachers, each specialising in a particular feature extraction aspect, we enhance the overall mental health classification performance. Through experimental validation, we demonstrate the efficacy of our model in achieving improved performance. All relevant codes will be made available upon publication.
mental health classification, knowledge distillation, multimodal
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: 33rd ACM International Conference on Information and Knowledge Management; October 21-25, 2024; Boise, Idahoisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Decision support systemsccs: Computing methodologies Information extraction
@@ 2.1. Mental Health Classification
Recent studies in mental health classification from social media content have incorporated diverse social media components. These components encompass various elements, including historical posts, conversation trees, social and interaction graphs, user or post metadata information, and profile pictures or posted images (Cao et al., 2019, 2022; Lin et al., 2020; Sawhney et al., 2021a, b). However, those additional sources will not always be available in the dataset due to data privacy restrictions or user preferences. This complicates research reproducibility since each study selects features based on what social media components are available to them. Hence, our research focuses on exploring mental health detection by analysing only social media textual posts, which is a compulsory component of text-based social media posts related to mental health. Based on the textual aspect, existing studies have worked on frequency- or score-based emotion features (Aragón et al., 2023; Zogan et al., 2022). More recent works use the fine-tuned contextual embeddings on emotion-based tasks as the emotion features for mental health classification (Lara et al., 2021; Sawhney et al., 2021a). These studies mainly focus on identifying and fixedly matching one type of emotion to each word or entire textual content. On the other hand, our research highlights the complexity of human emotions wherein a single word could be associated with multiple types of emotions by integrating the emotion-enriched features generated through a multi-label, corpus-based representation learning framework. In addition, we propose a new way to include acoustic features generated by original textual posts in this task, motivated by the proven effectiveness of vocal biomarkers in psychological distress and other medical conditions indication (Iyer et al., 2022). This would simultaneously process the emotion-enriched features as additional information. To the best of our knowledge, there have been no attempts to create a new modality from the text-only input to achieve unimodal text-based mental health risk detection tasks.
@@ 2.2. Multi-teacher Knowledge Distillation
To integrate the multimodal knowledge efficiently, we design our model with a knowledge distillation idea to compress a complex and large multimodal integration model into a smaller and simpler one while still retaining the accuracy and performance of the resultant model. Knowledge distillation (Hinton et al., 2015) involves transferring knowledge from a teacher model to a student model, commonly applied to compress large models by mapping intermediate layer outputs (Chen et al., 2021; Jiao et al., 2020) or minimising KL-divergence in class distribution (Mirzadeh et al., 2020). Traditionally, knowledge distillation is used within the same modality. However, recent approaches extend it to different modalities (Yang and Xu, 2021; Ni et al., 2022; Li et al., 2020). Some studies explore collaborative learning with multiple teachers for improved compression, such as (Wu et al., 2021) and (Pham et al., 2023) in language and vision models. (Tan et al., 2018) focuses on multilingual language translation, and (Vongkulbhisal et al., 2019) applies multi-teacher knowledge distillation to unify classifiers trained on distinct data sources. Inspired by this, we propose a new Multimodal Multi-Teacher Knowledge Distillation framework for mental health risk detection.
Figure 1. The overall architecture of 3M-Health: Multimodal Multi-teacher Knowledge Distillation for Mental Health Detection. Note that the teacher models are finetuned using each mental health dataset.
@@
In this section, we introduce the details of our Multimodal Multi-teacher knowledge distillation model for Mental Health detection, 3M-Health. Figure 1 illustrates the overall architecture design. This model consists of three distinct teacher models, each focusing on different modalities to independently learn diverse aspects of features crucial for interpreting mental health-related posts. The acquired features from multimodal teachers serve as a valuable source of knowledge for instructing the student model by utilising the average output distribution of the teacher models as soft targets. Note that we introduce three essential multimodal teacher models for mental health risk detection, including 1) a text-based teacher for understanding semantic aspects from input texts, 2) an emotion-based teacher for interpreting emotion aspects from input texts, and 3) an audio-based teacher for discerning emotions conveyed through audio sounds.
@@ 3.1.3. Audio-based Teacher
Existing social media mental health datasets primarily consist of text-only samples, so the two teachers mentioned before are purely based on the text. To address the need for a more comprehensive understanding of complex mental health and emotional contexts, we propose the integration of multimodality information. According to research (Collins et al., 2023), individuals can more accurately interpret the emotions of others through listening rather than observing facial expressions/body language or reading written text. Drawing inspiration from this insight, we introduce an audio-based teacher to enhance knowledge distillation, enabling the interpretation of emotions in mental health posts through sound. To achieve this, we first employ Bark [^4] https://github.com/suno-ai/bark, a pre-trained text-to-audio model, to generate corresponding audio for each post as Bark can capture emotional sounds detected from the text (e.g. [laughs], [gasps] and “…” for hesitations) [^5] A theoretical and practical comparison for text-to-audio generation APIs and a list of Bark’s sound cues in Section 4.2.. Note that Bark can generate only 13 seconds of audio. Hence, we tokenise each post at the sentence level, generate audio for each sentence, and then aggregate these audio segments into a complete audio representation for the entire post. Particularly long sentences or texts that do not have punctuation are further split into a maximum of 45 tokens.
@@ 3.2. Multimodal Multi-Teacher Fine-tuning
We performed fine-tuning on the text-based teacher, utilising the pre-trained language model, for the mental health classification task with labels $C=\{c_{1},c_{2},...,c_{|C|}\}$. This process enables the teacher to learn and adapt to the nuances of mental health-related contexts within each dataset.
For the emotion-based teacher, following the generation of emotion-rich representations for each post and its words, these embeddings serve as inputs for fine-tuning a Multi-Layer Perceptron (MLP) for mental health classification, operating over the labels $C$.
We employ the Audio Spectrogram Transformer (Gong et al., 2021) (AST) for the audio-based teacher to classify each generated audio into mental health risk classes. AST is a transformer-based model that takes a sequence of audio spectrogram patches as inputs. An audio waveform is first converted into a 128x100t spectrogram based on a sequence of 128-dimensional log Mel filterbank features computed with a 25ms Hamming window every 10ms. Such a spectrogram is then split into a sequence of N 16x16 patches of images with an overlap of 6 in both time and frequency dimensions. A special token [CLS] is added to the beginning of the sequence of spectrogram patches. After passing through transformer encoder layers, the [CLS] embedding is fed into a linear layer with sigmoid activation to classify mental health risk class labels $C$.
@@
3.3. Multi-Teacher Knowledge Distillation
4. Experimental Setup
4.1. Datasets
@@
We evaluate our proposed model using four publicly available datasets related to mental health on social media. Table 1 and Figure 2 provide a summary of statistics and an illustration of class distribution.
The TwitSuicide Dataset [^6] Data available upon request. (Long et al., 2022) replicates the data collection, processing, and annotation methods of (O’Dea et al., 2015). A sample of 660 tweets is annotated into three risk levels. The Strongly Concerning (SC) class is assigned to posts with a convincing display of severe suicidal ideation, while Safe to Ignore (SI) shows no evidence of suicide risk. If it doesn’t fall into other categories, a post remains in the Possibly Concerning (PC) class.
DEPTWEET [^7] https://github.com/mohsinulkabir14/DEPTWEET (Kabir et al., 2023) is collected from Twitter using seed terms based on the Patient Health Questionnaire (PHQ-9). The dataset comprises 40,191 tweets; however, only 5,128 tweets were retrieved during this study. The labels include Non-Depressed (ND), Mild (MI), Moderate (MO), and Severe (SE), maintaining an imbalanced class distribution, with around 80% labelled as ND and less than 2% SE.
The Identifying Depression Dataset [^8] https://github.com/Inusette/Identifying-depression (IdenDep) (Pirina and Çöltekin, 2018) consists of 1,841 Reddit posts, with “depression indicative” (DE) posts sourced from the Depression subreddit and non-depressive (NDE) posts from the “family” and “friendship advice” subreddits. No further manual check was done on the samples, increasing the probability of false negatives.
The SDCNL Dataset [^9] https://github.com/ayaanzhaque/SDCNL (Haque et al., 2021) involves distinguishing between Reddit suicide-related and depression-related posts. The dataset contains 1,895 nearly balanced posts labelled as Suicide (SUI) or Depression/Not Suicide (DEP) based on their subreddit. In accordance with (Benton et al., 2017), all posts are de-identified before any analysis, audio generation, and model training.
@@
We provide a detailed breakdown of text and audio statistics for each dataset class in Tables 2 and 3 to provide more information regarding the nature of each class, which may influence model learning and performance. Notably, DEPTWEET and IdenDep datasets have highly skewed data, with 82.16% and 70.23% on a single class, respectively. Figure 3 illustrates the differences in the generated audio in terms of duration. The Reddit-based datasets, IdenDep and SDCNL, are significantly longer than the Twitter-based datasets, possibly providing more auditory information inferred from the textual posts.
@@ 4.2. Text-to-Audio Generators
  1. (1)
    Tacotron2101010https://github.com/NVIDIA/tacotron2 (Shen et al., 2018) uses a recurrent neural network architecture to predict mel spectrogram sequences from text followed by a modified WaveNet vocoder.
  2. (2)
    SpeechT5111111https://github.com/microsoft/SpeechT5 (Ao et al., 2022) unifies modalities with a shared encoder-decoder architecture that uses cross-modal vector quantisation for speech and text alignment.
  3. (3)
    SpeechBrain [^12] https://github.com/speechbrain/speechbrain/ (Ravanelli et al., 2021) is a speech toolkit offering various speech related tasks. Their text-to-speech model is based on Tacotron2 but is trained further on the LJSpeech (Ito and Johnson, 2017) and LibriTTS (Zen et al., 2019) datasets.
  4. (4)
    Balacoon [^13] https://huggingface.co/balacoon/tts packages offer lightweight and fast text analysis and speech generation going against larger but slower TTS models. It sacrifices multi-speaker and multi-lingual features for lightning-fast speed on the CPU. The detailed model architecture was not publicly available at the time of this paper’s writing.
  5. (5)
    Bark [^14] https://github.com/suno-ai/bark has a GPT-based architecture using a quantised audio representation that does not require the use of phonemes allowing it to generalise beyond speech, thus making it a text-to-audio model.
@@ 4.4. Implementation Details
Hyperparameter tuning is done per dataset and model setup using Optuna [^16] https://optuna.org/ for 50 trials optimising weighted F1 scores. Detailed search space and best-found hyperparameters may be found in the Supplementary Material. Text-based teachers are trained using ReLu and a max length of 256. Audio-based teachers are trained for 25 epochs with a 5-epoch early stop, a 512 max length, and using the ReduceLRonPlateau scheduler. All inputs for the AST model are normalised to zero mean and 0.5 standard deviation. Emotion-based teachers are trained for 100 epochs with a 10-epoch early stop and a 256 max length. The student models are tuned and trained with distilled knowledge from the fine-tuned teachers using a max length of 256. All tuning was done using a 90:10 validation split and was conducted separately from the final model construction. All models are trained using an Adam optimiser on an NVIDIA TITAN RTX machine.
@@ 5.2. Audio Representation Analysis
1470 NDE *** introduce my girlfriend [18F] to my family My girlfriend lives in *** I live in *** *** introduce her to my mom but I don’t know how
@@ 5.3. Effectiveness of Multimodal Multi-Teachers
Figure 5. Distribution of multi-label emotion class labels for each dataset.
In general, the multi-teacher structure outperforms the use of a singular text-based teacher, although the effectiveness of different modalities varies across datasets. For the Twitter-based datasets, applying the emotion-based and text-based teachers together achieves the best results. In contrast, for Reddit-based datasets, the audio- and text-based teachers have better performances. This difference may be attributed to the longer posts in Reddit-based datasets, resulting in longer audio (Table 1) that contains more acoustic information beneficial for the audio-based teacher.
Moreover, due to the lengthier nature of posts in Reddit-based datasets, more SenticNet lexicon tokens are likely to be matched compared to Twitter-based datasets. This results in a higher number of generated emotion labels during the learning process of the multi-label emotion-based teacher. In Figure 5, we compare the count of generated multi-label emotion classes utilised to train the emotion-based teacher across all four datasets. It is evident that a greater proportion of posts in Reddit-based datasets match all seven emotion labels (shown in Section 3.1.2) compared to Twitter-based datasets. This potential increase in the number of matching emotion labels may present challenges in distinguishing between different emotions during the training of the emotion-based teacher, potentially impacting downstream mental health classification, especially for ambiguous classes such as Suicide (SUI) and Depression (DEP) in the SDCNL dataset.
5.4. Impact of Text-based Teachers
@@
We compare the impact of various pre-trained language models on the effectiveness of the text-based teacher. According to the results in Table LABEL:tbl:ablationstudytextteachers, BERT performs the best in terms of the weighted F1 across all four datasets. However, the domain-specific PLMs, such as MentalBERT or ClinicalBERT, perform better for most of the more concerning classes. For instance, ClinicalBERT outperforms BERT by 6.66% in the F1 score for the Strongly Concerning (SC) class in the TwitSuicide dataset, and MentalBERT achieves a 41.38% F1 score for the Severe (SE) class in the DEPTWEET dataset, surpassing all other language models that fail to recognise it. To ensure optimal performance of the text-based teacher, we specifically employ MentalBERT for experiments on the DEPTWEET dataset, while BERT is used for the other three datasets. Nonetheless, the overall performance of the text-based teacher is not significantly impacted by the choice of pre-trained language models. Notably, more performance enhancement stems from the inclusion of different modalities, as mentioned in the previous sections.

### v3

@@ Abstract.
The significance of mental health classification is paramount in contemporary society, where digital platforms serve as crucial sources for monitoring individuals’ well-being. However, existing social media mental health datasets primarily consist of text-only samples, potentially limiting the efficacy of models trained on such data. Recognising that humans utilise cross-modal information to comprehend complex situations or issues, we present a novel approach to address the limitations of current methodologies. In this work, we introduce a Multimodal and Multi-Teacher Knowledge Distillation model for Mental Health Classification, leveraging insights from cross-modal human understanding. Unlike conventional approaches that often rely on simple concatenation to integrate diverse features, our model addresses the challenge of appropriately representing inputs of varying natures (e.g., texts and sounds). To mitigate the computational complexity associated with integrating all features into a single model, we employ a multimodal and multi-teacher architecture. By distributing the learning process across multiple teachers, each specialising in a particular feature extraction aspect, we enhance the overall mental health classification performance. Through experimental validation, we demonstrate the efficacy of our model in achieving improved performance. [^1] Code available at https://github.com/adlnlp/3mhealth
mental health classification; knowledge distillation; multimodal
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679635isbn: 979-8-4007-0436-9/24/10ccs: Information systems Decision support systemsccs: Computing methodologies Information extraction
@@ 2.1. Mental Health Classification
Recent studies in mental health classification from social media content have incorporated diverse social media components. These components encompass various elements, including historical posts, conversation trees, social and interaction graphs, user or post metadata information, and profile pictures or posted images (Cao et al., 2019, 2022; Lin et al., 2020; Sawhney et al., 2021a, b). However, those additional sources will not always be available in the dataset due to data privacy restrictions or user preferences. This complicates research reproducibility since each study selects features based on what social media components are available to them. Our research focuses on exploring mental health detection by analysing only social media textual posts, which is a compulsory component of text-based social media posts related to mental health. Based on the textual aspect, existing studies have worked on frequency- or score-based emotion features (Aragón et al., 2023; Zogan et al., 2022). More recent works fine-tuned contextual embeddings on emotion-based tasks to use as emotion features (Lara et al., 2021; Sawhney et al., 2021a). These studies mainly focus on identifying and matching one type of emotion to each word or entire textual content. On the other hand, our research highlights the complexity of human emotions wherein a single word could be associated with multiple types of emotions by integrating the emotion-enriched features generated through a multi-label, corpus-based representation learning framework. In addition, we propose a new way to include acoustic features generated by original textual posts in this task, motivated by the effectiveness of vocal biomarkers in psychological distress and other medical condition indications (Iyer et al., 2022). This simultaneously processes the emotion-enriched features as additional information. To the best of our knowledge, there have been no attempts to create a new modality from the text-only input to achieve unimodal text-based mental health risk detection tasks.
@@ 2.2. Multi-teacher Knowledge Distillation
Other multimodal social media-based mental health detection studies mainly integrate modalities through concatenation of features (Han et al., 2020; Bucur et al., 2023) or a joint encoder (Anshul et al., 2024) which doesn’t represent the modalities properly. To integrate the multimodal knowledge efficiently, we design our model using knowledge distillation to compress a complex and large multimodal integration model into a smaller and simpler one while still retaining the accuracy and performance of the resultant model. Knowledge distillation (Hinton et al., 2015) involves transferring knowledge from a teacher model to a student model, commonly applied to compress large models by mapping intermediate layer outputs (Chen et al., 2021; Jiao et al., 2020) or minimising KL-divergence in class distribution (Mirzadeh et al., 2020). Traditionally, knowledge distillation is used within the same modality. However, recent approaches extend it to different modalities (Yang and Xu, 2021; Ni et al., 2022; Li et al., 2020). Some studies explore collaborative learning with multiple teachers for improved compression such as (Wu et al., 2021) and (Pham et al., 2023) in language and vision models. (Tan et al., 2018) focuses on multilingual language translation, and (Vongkulbhisal et al., 2019) applies multi-teacher knowledge distillation to unify classifiers trained on distinct data sources. Inspired by this, we propose a new Multimodal Multi-Teacher Knowledge Distillation framework for mental health risk detection.
Figure 1. Architecture of 3M-Health: Multimodal Multi-teacher Knowledge Distillation for Mental Health Detection.
@@
In this section, we introduce our Multimodal Multi-teacher knowledge distillation model for Mental Health detection, 3M-Health. Figure 1 illustrates the overall architecture. This model consists of three distinct teacher models, each focusing on different modalities to independently learn diverse aspects of features crucial for interpreting mental health-related posts. The acquired features from multimodal teachers serve as a valuable source of knowledge for instructing the student model by utilising the average output distribution of the teacher models as soft targets. We introduce three essential multimodal teacher models for mental health risk detection, including 1) a text-based teacher for understanding semantic aspects from input texts, 2) an emotion-based teacher for interpreting emotion aspects from input texts, and 3) an audio-based teacher for discerning emotions conveyed through audio sounds.
@@ 3.1.3. Audio-based Teacher
Existing social media mental health datasets primarily consist of text-only samples, so the two teachers mentioned before are purely based on the text. To address the need for a more comprehensive understanding of complex mental health and emotional contexts, we propose the integration of multimodality information. According to research (Collins et al., 2023), individuals can more accurately interpret the emotions of others through listening rather than observing facial expressions/body language or reading written text. Drawing inspiration from this insight, we introduce an audio-based teacher to enhance knowledge distillation, enabling the interpretation of emotions in mental health posts through sound. To achieve this, we first employ Bark, a pre-trained text-to-audio model, to generate corresponding audio for each post as Bark can capture emotional sounds detected from the text (e.g. [laughs], [gasps] and “…” for hesitations) [^5] A theoretical and practical comparison of text-to-audio generation APIs and a list of Bark’s sound cues in Section 4.2.. Note that Bark can generate only 13 seconds of audio. Hence, we tokenise each post at the sentence level, generate audio for each sentence, and then aggregate these audio segments into a complete audio representation for the entire post. Particularly long sentences or texts that do not have punctuation are further split into a maximum of 45 tokens.
@@ 3.2. Multimodal Multi-Teacher Fine-tuning
We fine-tuned pre-trained language models for the mental health classification task with labels $C=\{c_{1},c_{2},...,c_{|C|}\}$ to build the text-based teachers. This process enables the teacher to learn the nuances of mental health-related contexts within each dataset. For the emotion-based teacher, following the generation of emotion-rich representations for each post and its words, these embeddings serve as inputs for fine-tuning a Multi-Layer Perceptron (MLP) for mental health classification, operating over the labels $C$. We employ the Audio Spectrogram Transformer (Gong et al., 2021) (AST) for the audio-based teacher to classify each generated audio into mental health risk classes. AST is a transformer-based model that takes a sequence of audio spectrogram patches as inputs. An audio waveform is first converted into a 128x100t spectrogram based on a sequence of 128-dimensional log Mel filterbank features computed with a 25ms Hamming window every 10ms. Such a spectrogram is then split into a sequence of N 16x16 patches of images with an overlap of 6 in both time and frequency dimensions. A special token [CLS] is added to the beginning of the sequence of spectrogram patches. After passing through transformer encoder layers, the [CLS] embedding is fed into a linear layer with sigmoid activation to classify mental health risk class labels $C$.
@@
3.3. Multi-Teacher Knowledge Distillation
4. Experimental Setup
4.1. Datasets
@@
We evaluate our proposed model using four publicly available datasets related to mental health on social media. Table 1 and Figure 2 provide a summary of statistics and class distribution.
The TwitSuicide Dataset [^6] Data available upon request. (Long et al., 2022) replicates the data collection, processing, and annotation methods of (O’Dea et al., 2015). A sample of 660 tweets is annotated into three risk levels. The Strongly Concerning (SC) class is assigned to posts with a convincing display of severe suicidal ideation, while Safe to Ignore (SI) shows no evidence of suicide risk. If it doesn’t fall into other categories, a post remains in the Possibly Concerning (PC) class. DEPTWEET [^7] https://github.com/mohsinulkabir14/DEPTWEET (Kabir et al., 2023) is collected from Twitter using seed terms based on the Patient Health Questionnaire (PHQ-9). The dataset comprises 40,191 tweets; however, only 5,128 tweets were retrieved during this study. The labels include Non-Depressed (ND), Mild (MI), Moderate (MO), and Severe (SE), maintaining an imbalanced class distribution, with around 80% labelled as ND and less than 2% SE. The Identifying Depression Dataset [^8] https://github.com/Inusette/Identifying-depression (IdenDep) (Pirina and Çöltekin, 2018) consists of 1,841 Reddit posts, with “depression indicative” (DE) posts sourced from the Depression subreddit and non-depressive (NDE) posts from the “family” and “friendship advice” subreddits. No further manual check was done on the samples, increasing the probability of false negatives. The SDCNL Dataset [^9] https://github.com/ayaanzhaque/SDCNL (Haque et al., 2021) involves distinguishing between Reddit suicide-related and depression-related posts. The dataset contains 1,895 nearly balanced posts labelled as Suicide (SUI) or Depression/Not Suicide (DEP) based on their subreddit. In accordance with (Benton et al., 2017), all posts are de-identified before any analysis, audio generation, and model training.
We provide a detailed breakdown of text and audio statistics in Tables 2 and 3 to provide more information regarding the nature of each class, which may influence model learning and performance. Notably, DEPTWEET and IdenDep datasets have highly skewed data, with 82.16% and 70.23% on a single class, respectively. Figure 3 illustrates the differences in the generated audio in terms of duration. The Reddit-based datasets, IdenDep and SDCNL, are significantly longer than the Twitter-based datasets, possibly providing more auditory information inferred from the textual posts.
@@ 4.2. Text-to-Audio Generators
Tacotron2101010https://github.com/NVIDIA/tacotron2 (Shen et al., 2018) uses a recurrent neural network architecture to predict mel spectrogram sequences from text followed by a modified WaveNet vocoder. SpeechT5111111https://github.com/microsoft/SpeechT5 (Ao et al., 2022) unifies modalities with a shared encoder-decoder architecture that uses cross-modal vector quantisation for speech and text alignment. SpeechBrain [^12] https://github.com/speechbrain/speechbrain/ (Ravanelli et al., 2021) is a speech toolkit offering various speech related tasks. Their text-to-speech model is based on Tacotron2 but is trained further on the LJSpeech (Ito and Johnson, 2017) and LibriTTS (Zen et al., 2019) datasets. Balacoon [^13] https://huggingface.co/balacoon/tts packages offer lightweight and fast text analysis and speech generation going against larger but slower TTS models. It sacrifices multi-speaker and multi-lingual features for lightning-fast speed on the CPU. The detailed model architecture was not publicly available at the time of this paper’s writing. Bark [^14] https://github.com/suno-ai/bark has a GPT-based architecture using a quantised audio representation that does not require the use of phonemes allowing it to generalise beyond speech, thus making it a text-to-audio model.
@@ 4.4. Implementation Details
Hyperparameter tuning is done per dataset and model setup using Optuna [^16] https://optuna.org/ optimising weighted F1 scores. Detailed search space and best-found hyperparameters may be found in the Supplementary Material. Text-based teachers are trained using ReLu and a max length of 256. Audio-based teachers are trained for 25 epochs with a 5-epoch early stop, a 512 max length, and using the ReduceLRonPlateau scheduler. All inputs for the AST model are normalised to zero mean and 0.5 standard deviation. Emotion-based teachers are trained for 100 epochs with a 10-epoch early stop and a 256 max length. The student models are tuned and trained with distilled knowledge from the fine-tuned teachers using a max length of 256. All tuning was done using a 90:10 split and was conducted separately from the final model construction. All models are trained using an Adam optimiser on an NVIDIA TITAN RTX machine.
@@ 5.2. Audio Representation Analysis
1470 NDE *** introduce my girlfriend [18F] to my family My girlfriend lives *** I live in *** *** introduce her to my mom but I don’t know how
@@
Figure 5. Distribution of multi-label emotion class labels.
@@ 5.3. Effectiveness of Multimodal Multi-Teachers
In general, the multi-teacher structure outperforms the use of a singular text-based teacher, although the effectiveness of different modalities varies across datasets. For the Twitter-based datasets, applying the emotion-based and text-based teachers together achieves the best results. In contrast, for Reddit-based datasets, the audio- and text-based teachers have better performances. This difference may be attributed to the longer posts in Reddit-based datasets, resulting in longer audio (Table 1) that contains more acoustic information beneficial for the audio-based teacher. Moreover, due to the lengthier nature of posts in Reddit-based datasets, more SenticNet lexicon tokens are likely to be matched compared to Twitter-based datasets. This results in a higher number of generated emotion labels during the learning process of the multi-label emotion-based teacher. In Figure 5, we compare the count of generated multi-label emotion classes utilised to train the emotion-based teacher across all four datasets. It is evident that a greater proportion of posts in Reddit-based datasets match all seven emotion labels (shown in Section 3.1.2) compared to Twitter-based datasets. This potential increase in the number of matching emotion labels may present challenges in distinguishing between different emotions during the training of the emotion-based teacher, potentially impacting downstream mental health classification, especially for ambiguous classes such as Suicide (SUI) and Depression (DEP) in the SDCNL dataset.
5.4. Impact of Text-based Teachers
We compare the effectiveness of various PLMs for the text-based teacher. Table LABEL:tbl:ablationstudytextteachers shows that BERT produces the best weighted F1 across all datasets. However, the domain-specific PLMs perform better for the more concerning classes in a multi-class setup. ClinicalBERT outperforms BERT by 6.66% for Strongly Concerning (SC) in the TwitSuicide dataset, while MentalBERT achieves 41.38% for Severe (SE) in the DEPTWEET dataset, surpassing the other language models which failed to recognise it. To ensure optimal performance, we specifically employ MentalBERT for DEPTWEET, while BERT is used for the other datasets. Nonetheless, the overall performance of the text-based teacher is not significantly impacted by the choice of PLMs. More performance enhancement stems from the inclusion of different modalities, as discussed in the previous sections.